OpenStudy (anonymous):

Factoring Trinomials 2m^2+23m+63

OpenStudy (anonymous):

\[=2m^2+23m+63\] \[=2m^2+14m+9m+63\] \[=2m(m+7)+9(m+7)\] \[=(m+7)(2m+9)\]

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

2*63=126 14+9=23 14*9=126